# A Bayesian approach for fitting semi-Markov mixture models of cancer latency to individual-level data

### v1

@@ Abstract
Joint senior authors
Multi-state models of cancer natural history are widely used for designing and evaluating cancer early detection strategies. Calibrating such models against longitudinal data from screened cohorts is challenging, especially when fitting non-Markovian mixture models against individual-level data. Here, we consider a family of semi-Markov mixture models of cancer natural history introduce an efficient data-augmented Markov chain Monte Carlo sampling algorithm for fitting these models to individual-level screening and cancer diagnosis histories. Our fully Bayesian approach supports rigorous uncertainty quantification and model selection through leave-one-out cross-validation, and it enables the estimation of screening-related overdiagnosis rates. We demonstrate the effectiveness of our approach using synthetic data, showing that the sampling algorithm efficiently explores the joint posterior distribution of model parameters and latent variables. Finally, we apply our method to data from the US Breast Cancer Surveillance Consortium and estimate the extent of breast cancer overdiagnosis associated with mammography screening. The sampler and model comparison method are available in the R package baclava.
Keywords Cancer natural history  $\cdot$ Cancer indolence  $\cdot$ Overdiagnosis  $\cdot$ Data augmentation  $\cdot$ Markov chain Monte Carlo
@@ 1 Introduction
Cancer screening aims to detect early signs or markers of cancer before symptoms arise, facilitating timely intervention for improved patient outcomes. The efficacy of screening depends on the duration of the pre-clinical phase, during which the tumor is asymptomatic but screen-detectable. If the latency period of a tumor is too short, even frequent screening does not significantly advance the time of diagnosis. Conversely, a long latency period can result in overdiagnosis—the diagnosis and unnecessary treatment of slowly growing tumors that would not have caused symptoms or other harm in the person’s remaining lifetime (Welch and Black, 2010; Duffy and Parmar, 2013). Understanding the duration of tumor latency is thus crucial for assessing the balance between the benefits and harms of a screening test.
@@
Estimating tumor latency from cancer incidence data is challenging, because the onset of pre-clinical disease in healthy individuals is not directly observable. In unscreened populations, only the time of clinical diagnosis, which coincides with the end of the tumor latency period, is known. In this case, we cannot distinguish the time intervals before and after pre-clinical onset. In other words, the period of tumor latency is not identifiable. Fortunately, screening methods that utilize biological or imaging markers offer a selective insight into the latent disease state, thus enabling the identification of tumor latency.
@@
There is a rich history on estimating cancer latency from screening data using multi-state models that dates back to the 1980s. These models range from simple three-state models (Day and Walter, 1984; Shen and Zelen, 1999)—in which individuals are healthy (disease-free), pre-clinical (with an asymptomatic screen-detectable cancer) or clinical (with a symptomatic cancer)—to more complex models that seek to emulate the biological complexity of cancer progression (Rutter and Savarino, 2010; Seigneurin et al., 2011). Typically, the underlying stochastic processes of cancer natural history models are assumed to be Markovian for mathematical tractability. Under this assumption, the unobserved onset times can be integrated out analytically, giving a closed-form expression for the likelihood of observed data. This enables maximum likelihood estimation (Olsen et al., 2006; Wu et al., 2018) and Bayesian inference (Duffy et al., 2005; Ryser et al., 2022). However, relaxing the Markov assumption of state transitions often leads to a departure from the simpler mathematical structures, and the calibration of more complex non-Markovian models is usually performed using simulation-based or “likelihood-free” approximation methods, such as approximate Bayesian computation (Rutter et al., 2009; Alagoz et al., 2018; Bondi et al., 2023) or pseudo-likelihood methods (Moore et al., 2001).
@@
In this paper, we present a method for exact Bayesian inference of cancer natural history using a flexible family of semi-Markov mixture models that account for both progressive and non-progressive (indolent) disease states. The proposed approach accommodates individual-level screening and cancer diagnosis histories, and, because it does not rely on model-based simulation, is scalable to large cohorts of screened individuals. By relaxing the Markov assumption, we may then flexibly specify the hazard function, which in turn leads to improved model realism (Hsieh et al., 2002; Cheung et al., 2022). Since the semi-Markov models do not admit a closed form expression for the observed data likelihood, we conduct inference using data augmentation (Tanner and Wong, 1987). This strategy augments the observed data with latent variables in such a way that the resulting complete-data likelihood is amenable to iterative sampling. Executing this idea in the context of cancer natural history modeling is nontrivial because there are twice as many latent variables as the number of individuals in the study, and cancer screening cohorts often comprise tens or hundreds of thousands of individuals. The resulting joint posterior distribution is complex and high-dimensional, rendering standard rejection or importance sampling approaches impractical. Moreover, gradient-based approaches, such as Hamiltonian Monte Carlo, which are well-suited for high-dimensional problems (Girolami and Calderhead, 2011), are not applicable to our family of models due to the presence of discrete latent variables, and of discontinuities in the likelihood from the discrete screening outcomes.
@@
Instead, our data-augmented Markov chain Monte Carlo (DA-MCMC) sampler efficiently explores the joint posterior distribution of the model parameters together with the latent variables, enjoying sound theoretical guarantees. This fully Bayesian approach automatically accounts for the uncertainty in the exact onset time of pre-clinical cancers, and allows the estimation of any relevant function of the parameters such as the overdiagnosis rate through their posterior distribution. In addition, because the complete-data likelihood has a simple closed form, we can rigorously compare and select models using an extension of approximate leave-one-out cross-validation (Gelfand et al., 1992; Gelfand, 1995; Vehtari et al., 2017) to latent data models.
@@
The paper is structured as follows. In Section 2, we introduce the underlying family of semi-Markov mixture models of cancer natural history; in Section 3, we describe the inferential framework; and, in Section 4, we detail the construction of the DA-MCMC algorithm. We examine the mixing properties of the sampler through simulation studies in Section 5 and apply it to estimate the extent of overdiagnosis in a real-world cohort of women undergoing mammography screening in Section 6. The DA-MCMC sampling algorithm is publicly available in the R package baclava, and R code for reproducing the results present in this paper is available on GitHub (https://github.com/rmorsomme/baclava-manuscript).
@@ 2 Semi-Markov mixture model of cancer progression
(a) Compartmental model for cancer progression with a mixture of progressive and indolent pre-clinical cancers.
@@
(b) Illustration of possible individual clinical trajectories through the compartmental model.
Figure 1: Panel (a) depicts the compartmental model for the progression of cancer. Panel (b) illustrates possible individual trajectories through the compartmental model. Individuals start in compartment $H$. Only transitions from $H$ to $P$ and from $P$ to $C$ are possible, with compartment $P$ being an absorbing state for indolent individuals and compartment $C$ being absorbing for all others. In Panel (b), individual 1 does not develop pre-clinical cancer before the end of the observation period and, therefore, remains in $H$ until that time. Individual 2 develops pre-clinical cancer at age $\tau^{HP}_{2}$, at which time they transition to $P$, but they do not develop clinical cancer before the end of the observation period and, therefore, remain in $P$. Note that the pre-clinical cancer of individual 2 may be either indolent, in which case $\tau^{PC}_{2}=\infty$, or progressive with a clinical onset time after the end of the observation period. Individual 3 develops pre-clinical cancer at age $\tau^{HP}_{3}$ and transitions to state $P$. At age $\tau^{PC}_{3}$, they develop clinical cancer and thus transition to $C$. As $C$ is absorbing, the observation period ends with their transition time into $C$.
We model the dynamics of cancer progression as a multi-state mixture model. The model, depicted in Figure 1(a), consists of the three compartments healthy ($H$), pre-clinical cancer ($P$) and clinical cancer ($C$). Following Ryser et al. (2019), pre-clinical cancers are modeled as a mixture of progressive and indolent tumors. All individuals start in the healthy state $H$, and transition to a state of pre-clinical screen-detectable cancer $P$ after a randomly distributed time. Once in $P$, individuals with an indolent cancer stay in this compartment indefinitely, while individuals with a progressive cancer eventually transition to $C$ when their cancer becomes symptomatic.
Several early works consider purely progressive models, in which all pre-clinical tumors progress to clinical disease after a finite time (Day and Walter, 1984; Shen and Zelen, 1999, 2001). Allowing a proportion of pre-clinical cancers to be non-progressive is particularly important for slowly growing or indolent disease entities, such as low-risk prostate cancers or in situ breast cancers (Baker et al., 2014). Despite the challenges of fitting such mixture models in practice (Ryser et al., 2019), these models have gained in popularity, especially in the study of breast cancer progression (Chen et al., 1996; Duffy et al., 2005; Olsen et al., 2006; Seigneurin et al., 2011; Alagoz et al., 2018; Wu et al., 2018; Ryser et al., 2022).
@@
The associated stochastic process is characterized by the probabilistic mechanisms governing the transitions between the compartments. Denote the waiting time until pre-clinical cancer with $\sigma^{H}=\tau^{HP}-t_{0}$ and the pre-clinical sojourn time with $\sigma^{P}=\tau^{PC}-\tau^{HP}$, corresponding to the interval between the onset of pre-clinical cancer and the diagnosis with symptomatic disease. An individual with an indolent tumor has $\sigma^{P}=\infty$. We assume that
@@
independently, where $\textsf{F}_{\textsf{H}}$ is a distribution on the positive line, and $\textsf{F}_{\textsf{P}}$ corresponds to the mixture cure model (Berkson and Gage, 1952; Peng and Taylor, 2014),
@@
with a point mass at $\infty$ for indolent cancers and a density on the positive line for progressive cancers. Here, $\psi\in[0,1]$ is the proportion of indolent cancers and $\delta_{\infty}$ the Dirac distribution with point mass at $\infty$. Distribution (2) implies that a proportion $\psi$ of tumors are indolent with infinite sojourn times while the sojourn time of progressive tumors follows distribution $\textsf{F}_{\textsf{prog}}$.
Our inferential framework will apply to arbitrary choices of parametric distributions on the positive line for $\textsf{F}_{\textsf{H}}$ and $\textsf{F}_{\textsf{prog}}$, resulting in a flexible semi-Markovian formulation. We write $\theta$ to refer to the collective parameters of $\textsf{F}_{\textsf{H}}$ and $\textsf{F}_{\textsf{prog}}$. This formulation encompasses existing models as special cases: for example, letting $\textsf{F}_{\textsf{H}}$ and $\textsf{F}_{\textsf{prog}}$ be exponential distributions results in the Markov mixture model of Seigneurin et al. (2011), and $\psi=0$ recovers the purely progressive model of Shen and Zelen (1999, 2001). Introducing the indolence indicator $I$, which equals $1$ if the cancer is indolent and 00 otherwise, we can also express the distribution (2) hierarchically as
$I$ $\sim\textsf{Ber}(\psi),$ (3)
$\sigma^{P}\mid I$$\sim\begin{cases}\delta_{\infty},&\quad I=1\\ \textsf{F}_{\textsf{prog}},&\quad I=0\end{cases},$ (4)
@@ 3.1 Likelihood of complete data
Consider a sample of $n$ individuals that evolve independently, and let the subscript $i=1,\dots,n$ refer to quantities related to the $i$th individual. We refer to the data for individual $i$ as $X_{i}=\left(I_{i},\tau^{HP}_{i},\tau^{PC}_{i}\right)$ and write $\mathbf{X}=(X_{1},\dots,X_{n})$ for the data of the entire population. By independence between the individuals, the likelihood of the process described by equations (1)-(4) factorizes as
$\tilde{L}(\theta,\psi;\mathbf{X})=\prod_{i=1}^{n}\tilde{L}_{i}(\theta,\psi;X_{i})$ (5)
where
$\tilde{L}_{i}(\theta,\psi;X_{i})=\begin{cases}S_{\textsf{H}}\left(c_{i}-t_{0};\theta\right),&\quad c_{i}<\tau^{HP}_{i}\\ f_{\textsf{H}}\left(\sigma^{H}_{i};\theta\right)f_{\textsf{I}}(I_{i};\psi)S_{\textsf{prog}}\left(c_{i}-\tau^{HP}_{i};\theta\right)^{1-I_{i}},&\quad\tau^{HP}_{i}\leq c_{i}<\tau^{PC}_{i}\\ f_{\textsf{H}}\left(\sigma^{H}_{i};\theta\right)(1-\psi)f_{\textsf{prog}}\left(\sigma^{P}_{i};\theta\right),&\quad\tau^{PC}_{i}\leq c_{i}\end{cases}.$ (6)
@@ 3.2 Partially observed data
(b) Screen-detected individual.
@@
(c) Interval-detected individual
@@
$1$ $40$ 00
$1$ $45$ 00
$2$ $40$ 00
$2$ $45$ $1$
$3$ $40$ 00
$3$ $45$ 00
(d) Observed screen data for the 3 individuals corresponding to Figures 2(a)-2(c).
@@
$1$ $46.5$ $46.5$
$2$ $45$ $45$
$3$ $46$
(e) Observed right-censored transition age into $C$, $t^{PC}_{i}$, and censoring age for the 3 individuals corresponding to Figures 2(a)-2(c).
Figure 2: Observed data for the individuals depicted in Figure 1(b). Panels (a-c): diagram displaying the trajectories and screen results. Panel (d): screen results. Panel (e): right-censored onset age of clinical cancer. In Panel (a), the individual does not develop pre-clinical cancer before their censoring age. They are in $H$ at the two screen times, the two screens are therefore negative. They are censored at age $46.5$, $1.5$ years after their last screen. In Panel (b), the individual develops a pre-clinical before age $40$. They are in $P$ when the two screens take place. The first screen is a false negative, and the second is positive. They are censored at age $45$, their age at the positive screen. In Panel (c), the individual develops a pre-clinical cancer between the first and second screen and a clinical cancer at $46$. Their screen at age $45$ is a false negative. The onset time of their clinical cancer at age $46$ is observed.
We now turn our attention to real-world data sources, such as prospective screening trials and observational screening cohorts, In such settings, the data consist of a series of screens and associated results, as well as clinical cancer diagnoses. Let $n_{i}$ denote the number of screens available for person $i$, with their age and outcome at each of these screening times contained in the vectors $\left(t^{i}_{1},\dots t^{i}_{n_{i}}\right)$ and $\left(O^{i}_{1},\dots O^{i}_{n_{i}}\right)\in\{0,1\}^{n_{i}}$, respectively, where $t_{1}^{i}=e_{i}$ corresponds to the age at study entry. These data are complemented by the time $t^{PC}_{i}=\min\left\{c_{i},\tau^{PC}_{i}\right\}$, which corresponds to either the age of diagnosis with a screen-detected cancer, the age of diagnosis with a clinical cancer, or to the age of censoring otherwise.
We assume that the screening ages are non-informative, that is, that they are ignorable or independent of the disease process and of the model parameters. Under this assumption, we may condition on them in our analysis (Gelman et al., 2013). We therefore denote the observed variables of individual $i$ with $Y_{i}=\left(O^{i}_{1},\dots,O^{i}_{n_{i}},t^{PC}_{i}\right)$ and condition on the ignorable screen ages $(t^{i}_{1},\dots,t^{i}_{n_{i}})$. We further write $\mathbf{Y}=(Y_{1},\dots,Y_{n})$. Tables 2(d) and 2(e) present the observed data of a sample of three individuals for illustration.

### v2

@@ Abstract
[Abstract] Multi-state models of cancer natural history are widely used for designing and evaluating cancer early detection strategies. Calibrating such models against longitudinal data from screened cohorts is challenging, especially when fitting non-Markovian mixture models against individual-level data. Here, we consider a family of semi-Markov mixture models of cancer natural history and introduce an efficient data-augmented Markov chain Monte Carlo sampling algorithm for fitting these models to individual-level screening and cancer diagnosis histories. Our fully Bayesian approach supports rigorous uncertainty quantification and model selection through leave-one-out cross-validation, and it enables the estimation of screening-related overdiagnosis rates. We demonstrate the effectiveness of our approach using simulated data, showing that the sampling algorithm efficiently explores the joint posterior distribution of model parameters and latent variables. Finally, we apply our method to data from the US Breast Cancer Surveillance Consortium and estimate the extent of breast cancer overdiagnosis associated with mammography screening. The sampler and model comparison method are available in the R package baclava.
[…]
keywords:
cancer natural history, indolence, overdiagnosis, data augmentation, Markov chain Monte Carlo.
articletype: Article Typejournal: Statistics in Medicinevolume: 00footnotetext: Abbreviations: MCMC, Markov chain Monte Carlo; MH, Metropolis-Hastings; DA-MCMC, data-augmented Markov chain Monte Carlo; BCSC, Breast Cancer Surveillance Consortium; CV, cross-validation; ALOOCV, approximate leave-one-out cross-validation.
@@ 1 Introduction
Cancer screening aims to detect early markers of cancer before symptoms arise, facilitating timely intervention for improved patient outcomes. The efficacy of screening depends on the duration of the pre-clinical phase, during which a tumor is asymptomatic but screen-detectable. If the tumor latency period is too short, even frequent screening does not significantly advance the time of diagnosis. Conversely, a long latency period can result in overdiagnosis and overtreatment—the diagnosis and unnecessary treatment of tumors that would not have caused symptoms or other harm in the person’s remaining lifetime 1, 2. Understanding the duration of tumor latency is thus crucial for assessing the benefits and harms of interventions that follow positive screening tests.
@@
Estimating tumor latency from cancer incidence data is challenging because the onset of pre-clinical disease in healthy individuals is not directly observable. In unscreened populations, only the time of clinical diagnosis, which coincides with the end of the tumor latency period, is known. In this case, we cannot distinguish the time intervals before and after pre-clinical onset. In other words, the period of tumor latency is not identifiable. Fortunately, screening methods that utilize biological or imaging markers offer a selective insight into the latent disease state, thus enabling the identification of tumor latency.
@@
There is a rich history on estimating cancer latency from screening data using multi-state models that dates back to the 1980s. These models range from simple three-state models 3, 4—in which individuals are healthy (disease-free), pre-clinical (with an asymptomatic screen-detectable cancer) or clinical (with a symptomatic cancer)—to more complex models that seek to emulate the biological complexity of cancer progression 5, 6. Typically, the underlying stochastic processes of such multi-state models are assumed to be Markovian for mathematical tractability 7. Under this assumption, the unobserved onset times can be integrated out analytically, giving a closed-form expression for the likelihood of observed data. This enables maximum likelihood estimation 8, 9 and Bayesian inference 10, 11. However, relaxing the Markov assumption often leads to a departure from the simpler mathematical structures, and more complex non-Markovian models are usually calibrated using simulation-based or “likelihood-free” approximation methods, such as approximate Bayesian computation 12, 13, 14 or pseudo-likelihood methods 15.
@@
In this paper, we present a method for exact Bayesian inference of cancer natural history using a flexible family of semi-Markov mixture models that account for both progressive and non-progressive (indolent) disease states. The proposed approach accommodates individual-level screening and cancer diagnosis histories, and, because it does not rely on model-based simulation, is scalable to large cohorts of screened individuals. By relaxing the Markov assumption, we may then flexibly specify the hazard function, which in turn leads to improved model realism 16, 17. Since the semi-Markov models do not admit a closed form expression for the observed data likelihood, we conduct inference using data augmentation 18. This strategy augments the observed data with latent variables in such a way that the resulting complete-data likelihood is amenable to iterative sampling. Executing this idea in the context of cancer natural history modeling is nontrivial because the number of latent variables is proportional to the number of individuals in the study, and cancer screening cohorts often comprise tens or hundreds of thousands of individuals. The resulting joint posterior distribution is complex and high-dimensional, rendering standard rejection or importance sampling approaches impractical. Moreover, gradient-based approaches, such as Hamiltonian Monte Carlo, which are well-suited for high-dimensional problems 19, are not applicable to our family of models due to the presence of discrete latent variables, and of discontinuities in the likelihood from the discrete screening outcomes.
@@
Instead, our DA-MCMC sampler efficiently explores the joint posterior distribution of the model parameters together with the latent variables, enjoying sound theoretical guarantees. This fully Bayesian approach automatically accounts for the uncertainty in the unobserved onset time of cancers, and allows the estimation of any relevant function of the parameters, e.g. the overdiagnosis rate, through its posterior distribution. Moreover, because the complete-data likelihood has a tractable closed form, we can rigorously compare models using an extension of ALOOCV 20, 21 to latent models.
@@
The paper is structured as follows. In Section 2, we introduce the underlying family of semi-Markov mixture models of cancer natural history; in Section 3, we describe the inferential framework; and, in Section 4, we detail the construction of the DA-MCMC algorithm. We examine the mixing properties of the sampler through simulation studies in Section 5 and apply it to estimate the extent of overdiagnosis in a real-world cohort of women undergoing mammography screening in Section 6. The DA-MCMC sampler is publicly available in the R package implementation baclava on CRAN, and R code for reproducing the results present in this paper is available on GitHub (https://github.com/rmorsomme/baclava-manuscript).
@@ 2 Semi-Markov mixture model of cancer progression
((a)) Compartmental model for cancer progression with a mixture of progressive and indolent pre-clinical cancers.
@@
((b)) Illustration of possible individual clinical trajectories through the compartmental model.
Figure 1: Panel (a) depicts the compartmental model for the progression of cancer. Panel (b) illustrates possible individual trajectories through the compartmental model. Individuals start in compartment $H$. Only transitions from $H$ to $P$ and from $P$ to $C$ are possible, with compartment $P$ being absorbing for indolent individuals and compartment $C$ being absorbing for all others. In Panel (b), individual 1 does not develop pre-clinical cancer before the end of the observation period and, therefore, remains in $H$ until that time. Individual 2 develops pre-clinical cancer at age $\tau^{HP}_{2}$, at which time they transition to $P$, but they do not develop clinical cancer before the end of the observation period and, therefore, remain in $P$. Note that the pre-clinical cancer of individual 2 may be either indolent, in which case $\tau^{PC}_{2}=\infty$, or progressive with a clinical onset time after the end of the observation period. Individual 3 develops pre-clinical cancer at age $\tau^{HP}_{3}$ and transitions to state $P$. At age $\tau^{PC}_{3}$, they develop clinical cancer and thus transition to $C$. As $C$ is absorbing, the observation period ends with their transition time into $C$.
We model the dynamics of cancer progression as a multi-state mixture model. The model, depicted in Figure 1(a), consists of the three compartments healthy ($H$), pre-clinical cancer ($P$) and clinical cancer ($C$). Following 22, pre-clinical cancers are modeled as a mixture of progressive and indolent tumors. All individuals start in the healthy state $H$, and transition to a state of pre-clinical screen-detectable cancer $P$ after a randomly distributed time. Once in $P$, individuals with an indolent cancer stay in this compartment indefinitely, while individuals with a progressive cancer transition to $C$ when their cancer becomes symptomatic.
Several early works consider purely progressive models, in which all pre-clinical tumors progress to clinical disease after a finite time 3, 4, 23. Allowing a proportion of pre-clinical cancers to be non-progressive is particularly important for slowly growing or indolent disease entities, such as low-risk prostate cancers or in situ breast cancers 24. Despite the challenges of fitting such mixture models in practice 22, they have gained in popularity, especially in the study of breast cancer progression 25, 10, 8, 6, 13, 9, 11.
@@
The associated stochastic process is characterized by the probabilistic mechanisms governing the transitions between the compartments. Denote the waiting time until pre-clinical cancer with $\sigma^{H}=\tau^{HP}-t_{0}$ and the pre-clinical sojourn time with $\sigma^{P}=\tau^{PC}-\tau^{HP}$, corresponding to the interval between the onset of pre-clinical cancer and the symptomatic diagnosis. An individual with an indolent tumor has $\sigma^{P}=\infty$. We assume that
@@
independently, where $\textsf{F}_{\textsf{H}}$ is a distribution on the positive line, and $\textsf{F}_{\textsf{P}}$ corresponds to the mixture cure model 26,
@@
with a point mass at $\infty$ for indolent cancers and a density on the positive line for progressive cancers. Here, $\psi\in[0,1]$ is the proportion of indolent cancers and $\delta_{\infty}$ the Dirac distribution with its mass at $\infty$. Distribution (2) implies that a proportion $\psi$ of tumors are indolent with infinite sojourn times while the sojourn time of progressive tumors follows distribution $\textsf{F}_{\textsf{prog}}$.
Our inferential framework will apply to arbitrary choices of parametric distributions on the positive line for $\textsf{F}_{\textsf{H}}$ and $\textsf{F}_{\textsf{prog}}$, resulting in a flexible semi-Markovian formulation. We write $\theta$ to refer to the collective parameters of $\textsf{F}_{\textsf{H}}$ and $\textsf{F}_{\textsf{prog}}$. This formulation encompasses existing models as special cases: for example, letting $\textsf{F}_{\textsf{H}}$ and $\textsf{F}_{\textsf{prog}}$ be exponential distributions results in the Markov mixture model of 6, and $\psi=0$ recovers the purely progressive model of 4, 23. Introducing the indolence indicator $I$, which equals $1$ if the cancer is indolent and 0 otherwise, we can also express the distribution (2) hierarchically as
$I\sim\textsf{Ber}(\psi),\quad\sigma^{P}\mid I\sim\begin{cases}\delta_{\infty},&\quad I=1\\ \textsf{F}_{\textsf{prog}},&\quad I=0\end{cases},$ (3)
@@ 3.1 Likelihood of complete data
Consider a sample of $n$ independent individuals, and let the subscript $i=1,\dots,n$ refer to variables of the $i$th individual. We refer to the data for individual $i$ as $X_{i}=\left(I_{i},\tau^{HP}_{i},\tau^{PC}_{i}\right)$ and write $\mathbf{X}=(X_{1},\dots,X_{n})$ for the data of the entire population. By independence between the individuals, the likelihood of the process described by equations (1)-(3) factorizes as
$\tilde{L}(\theta,\psi;\mathbf{X})=\prod_{i=1}^{n}\tilde{L}_{i}(\theta,\psi;X_{i}),\quad\text{where}\quad\tilde{L}_{i}(\theta,\psi;X_{i})=\begin{cases}S_{\textsf{H}}\left(c_{i}-t_{0};\theta\right),&\quad c_{i}<\tau^{HP}_{i}\\ f_{\textsf{H}}\left(\sigma^{H}_{i};\theta\right)f_{\textsf{I}}(I_{i};\psi)S_{\textsf{prog}}\left(c_{i}-\tau^{HP}_{i};\theta\right)^{1-I_{i}},&\quad\tau^{HP}_{i}\leq c_{i}<\tau^{PC}_{i}\\ f_{\textsf{H}}\left(\sigma^{H}_{i};\theta\right)(1-\psi)f_{\textsf{prog}}\left(\sigma^{P}_{i};\theta\right),&\quad\tau^{PC}_{i}\leq c_{i}\end{cases}.$ (4)
@@ 3.2 Partially observed data
((b)) Screen-detected individual.
@@
((c)) Interval-detected individual
@@
11 $40$ 0
11 $45$ 0
22 $40$ 0
22 $45$ 11
33 $40$ 0
33 $45$ 0
((d)) Observed screen data for the 3 individuals corresponding to Figures 2(a)-2(c).
@@ 33 $45$ 0
11 $46.5$ $46.5$
22 $45$ $45$
33 $46$
((e)) Observed right-censored transition age into $C$, $t^{PC}_{i}$, and censoring age for the 3 individuals corresponding to Figures 2(a)-2(c).
Figure 2: Observed data for the individuals depicted in Figure 1(b). Panels (a-c): diagrams displaying the trajectories and screen results. Panel (d): screen results. Panel (e): right-censored onset age of clinical cancer. In Panel (a), the individual is in $H$ at the two screen times, the two screens are therefore negative. The individual is censored at age $46.5$, $1.5$ years after their last screen. In Panel (b), the individual is in $P$ when the two screens take place. The first screen is a false negative, and the second is positive. The individual is censored at age $45$, their age at the positive screen. In Panel (c), the screen at age $45$ is a false negative. The onset time of the clinical cancer at age $46$ is observed.
We now turn our attention to real-world data sources, such as prospective screening trials and observational screening cohorts, In such settings, the data consist of a series of screens and associated results, as well as clinical cancer diagnoses. Let $n_{i}$ denote the number of screens available for person $i$, with their age and outcome at each of these screening times contained in the vectors $\left(t^{i}_{1},\dots t^{i}_{n_{i}}\right)$ and $\left(O^{i}_{1},\dots O^{i}_{n_{i}}\right)\in\{0,1\}^{n_{i}}$, respectively, where $t_{1}^{i}=e_{i}$ corresponds to the age at study entry; $O^{i}_{j}=1$ when the $j$th screen of participant $i$ is positive and $O^{i}_{j}=0$ when it is negative. These data are complemented by the time $t^{PC}_{i}=\min\left\{c_{i},\tau^{PC}_{i}\right\}$, which corresponds to either the age of diagnosis with a screen-detected cancer, the age of diagnosis with a clinical cancer, or to the age of censoring otherwise, and by the indicator $\delta_{i}=\delta_{\tau^{PC}_{i}}\left(t^{PC}_{i}\right)$, which takes on the values $1$ when $t^{PC}_{i}=\tau^{PC}_{i}$, that is, for individuals with a clinical cancer, and 0 otherwise.
We assume that the screening ages are non-informative, that is, that they are ignorable or independent of the disease process and of the model parameters. Under this assumption, we may condition on them in our analysis 27. We therefore denote the observed variables of individual $i$ with $Y_{i}=\left(O^{i}_{1},\dots,O^{i}_{n_{i}},t^{PC}_{i},\delta_{i}\right)$ and condition on the ignorable screen ages $(t^{i}_{1},\dots,t^{i}_{n_{i}})$. We further write $\mathbf{Y}=(Y_{1},\dots,Y_{n})$. Tables 2(d) and 2(e) present the observed data of a sample of three individuals for illustration.